OpenStudy (anonymous):

I asked.... 6-(x/3)+1>(2/3) I got a good answer of.... -x+21>2 19>x I still cant figure out how we got -x+21>2

OpenStudy (swag):

x<19

OpenStudy (anonymous):

I dont know the first few steps please to get to -x+21>2, what are the steps PRIOR to that?

OpenStudy (anonymous):

start with \[6-\frac{x}{3}+1>\frac{2}{3}\] combine like terms to get \[7-\frac{x}{3}>\frac{2}{3}\] now if the fractions annoy you multiply both sides by 3 to get \[21-x>2\]

OpenStudy (anonymous):

thank you